OpenStudy (anonymous):

A certain sum of money is invested at 10%. Twice that amount is invested at 8%. The total amount of interest from both investments is $91. how much is invested at 8%.

OpenStudy (amistre64):

Hmmm.... interest is Money times %. So if we say x(10%) + 2x(8%) = 91 we should be able to solve for x. x(.1) + 2x(.08) = 91 x(.1 + 2(.08)) = 91 x(.1 + .16) = 91 x(.26) = 91 x=350 and 2x = 700. Double check: 350(.1) + 700(.08) = 91 35 + 56 = 91 91 = 91 ... yay!!! 700 was invested at 8%. If I did it correctly that is :) x = 91/.26 = 9100/26 x=